# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v1

@@ Abstract
We propose to solve the SUSY $\mu$-problem in the framework of modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry and the effective $\mu$-term is re-generated after the modulus field is stabilized. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. The presence of $X_{nl}$ contributions can possibly slightly change the original stabilized modulus VEV to a value located in the vicinity of some fixed points. Natural value of $\mu_{eff}$, which is much smaller than the SUSY scale, can either be the consequence of expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), or be the joint consequence of suppression from both powers of $q$ [or $(2\Im\tau)^{-1}$] and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
@@ 1 Introduction
Low energy supersymmetry (SUSY) is well motivated theoretically and regarded as one of the most appealing extensions of the Standard Model. Not only the quadratic divergences related to the Higgs mass can be eliminated, but also successful gauge coupling unification can be realized in the framework of low energy SUSY. Especially, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ required by minimal supersymmetric standard model (MSSM) and it can be seen as a triumph of low energy SUSY.
Although the framework of low energy SUSY is attracting, it still has several unsatisfactory features. In addition to the problem of null observation of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem: the $\mu$-parameter is a SUSY-respecting parameter appearing in the superpotential, while it is required to be roughly of order the soft SUSY breaking scale to allow a Higgs VEV of order 174 GeV without too much miraculous cancelation.
@@
There are many proposals to tackle the $\mu$-problem, for example, by singlet extension to next-to-minimal supersymmetric standard model (NMSSM) or by non-renormalizable Lagrangian terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ). Most solutions consist of two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is forbidden, usually by some discrete or continuous symmetry; Second, the $\mu$ term is re-generated via some mechanism such as symmetry breaking. One can see in Bae:2019dgg some detailed realizations of the two steps for the $\mu$-problem.
@@
We would like to tackle the $\mu$-problem with a similar but different approach. The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry. Besides, the re-generation of $\mu$ term is realized after modular symmetry breaking with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions. Naturally suppressed $\mu$-parameter can thus be realized.
@@
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, (exact) moduli spaces of Calabi-Yau threefolds, like the quintic and fairly generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . The modular symmetry can constrain stringently the superpotential as the relevant Yukawa couplings are modular forms. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In modular flavor symmetry framework, the masses and mixings for the quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been adopted in various studies to explain the low energy flavor structures of the Standard Model. Besides, modular flavor symmetries can combine with the GUT framework to increase greatly the prediction ability of the models, such as the combination of $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
@@
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking, it can also be the only source of CP symmetry breaking. The VEV of the modulus field $\tau$ will in general break the modular symmetry completely except three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms, thus the Yukawa coupling, and the fermion mass matrices will be fixed after the modular flavour symmetry is fully or partially broken by VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with modular-invariant scalar potential in 4d $N=1$ supergravity, which is also crucial to give nonzero masses for modulus field to render it phenomenological acceptable.
@@
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induces a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane like sources can be described in terms of a nilpotent goldstino superfield KS in a supersymmetric effective field. Although SUSY is non-linearly realized, the couplings of chiral, gauge and gravity superfields coupled to the constrained nilpotent goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action.
@@
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem with modular symmetry. In Sec 4, we discuss the stabilization of singlet modulus field, with the presence of SUSY breaking contributions described in the non-linear realized SUSY scheme with nilpotent goldstino $X_{nl}$ superfield. In Sec 5, we discuss the re-generation of naturally small $\mu_{eff}$ after modulus stabilization. Sec 6 contains our conclusions.
@@ 2 Modular $\Gamma_{N}$ symmetry
The full modular group is the group of linear fraction transformations acting on the complex modulus $\tau$ in the upper half complex plane
@@
It can be generated by two generators $S$ and $T$
@@
which satisfy
@@
The two generators can be represented by the following $SL(2,Z)$ matrices,
@@
Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined by $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, with $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ defined by
$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.$ (7)
@@
Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension. They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{2k}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
@@ 3 Natural $\mu$-term in modular flavor model
As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of $\mu$-parameter in the superpotential should be forbidden by some symmetry. We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformation so as that ordinary $\mu H_{u}H_{d}$ term is forbidden by modular symmetry.
@@
with the coefficient for bilinear $H_{u}H_{d}$ term given by
$\mu=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ with $62k_{3}\in 24\mathbb{Z}+$ 6 to guarantee the modular invariance of
@@
Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale. The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase, whose transformation under $\gamma$ can be given as
$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{}.$ (15)
with $\epsilon_{\eta}(\gamma)$ depending on $\gamma$ (but not on $\tau$) and being given in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right].$ (16)
@@
It is obvious that the Kahler potential for superfields with non-trivial modular weights, for example, the $H_{u}$ and $H_{d}$ superfields, are non-minimal. Therefore, after the moduli fields $\tau$ acquire a VEV by modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical forms, which amounts to the replacements
$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{u}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{d}/2}~{}.$ (18)
So, when the modulus field is fixed by modulus stabilization, the effective $\mu$-parameter will be re-generated to be
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{u}+k_{d})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why such $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and be of order the soft SUSY breaking scale $m_{soft}$. We indicate that the smallness of $\mu_{eff}$ is the consequence of large suppression factor, either origin from some power of deviation of modulus VEV from typical fixed point or from the non-trivial power (that determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of some factor that related to the modulus VEV. All of the suppression factors are determined by the modular stabilization mechanism, with the stabilized modulus VEVs affected by SUSY breaking effects.
4 Stabilization of modulus field with SUSY breaking effects
To stabilize the modulus field, effective supergravity scalar potentials from weight $-3$ modular form superpotential can be used to determine the minimum. The modulus VEV can always be restricted to the fundamental domain $D_{H}$
${D}_{F}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{F}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice. Discussions on the modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
To guarantee the modular invariance and Kahler transformation invariance of

### v3

@@ Abstract
We propose a solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry, and an effective $\mu$-term is regenerated following the stabilization of the modulus field. We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield. The inclusion of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near some fixed point. A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
@@ 1 Introduction
Low-energy supersymmetry (SUSY) is well motivated from a theoretical standpoint and is considered to be one of the most compelling extensions of the Standard Model. It not only address the quadratic divergence problem associated with the Higgs mass but also enables successful gauge coupling unification. Notably, the observed 125 GeV scalar, discovered by both the ATLAS ATLAS and CMS CMS collaborations of LHC, falls miraculously within the narrow $115\sim 135$ GeV ’window’ predicted by the minimal supersymmetric standard model (MSSM). This can be regarded as strong supports for low energy SUSY.
@@
Although the framework of low energy SUSY is attractive, it still has several unsatisfactory features. In addition to the lack of conclusive signals for superparticles at the LHC, the MSSM is also bothered by the SUSY $\mu$-problem. The $\mu$-parameter, a SUSY preserving parameter in the superpotential, however, is required (without too much miraculous cancelation) to be roughly of the order of the soft SUSY breaking scale, as determined by the minimization of the MSSM Higgs potential. Many proposals have been put forward to address the $\mu$-problem, for example, by singlet extension of MSSM to next-to-minimal supersymmetric standard model (NMSSM) or by introducing non-renormalizable terms in the superpotential via Kim-Nilles mechanism KN (or in the Kahler potential via Giudice-Masiero mechanism GM ). Most solutions involve two steps Bae:2019dgg : First, the appearance of $\mu$-parameter is prohibited, usually by some discrete or continuous symmetry. Second, the $\mu$ term is regenerated via some mechanism such as symmetry breaking. For detailed realizations of these two steps, one can refer to Bae:2019dgg .
@@
We propose to tackle the $\mu$-problem with a similar but different approach. The symmetry adopted to forbid the $\mu$-parameter is a modular flavor symmetry. The regeneration of $\mu$ term is realized following the modular symmetry breaking, with the modulus stabilization mechanism accompanied by typical SUSY breaking contributions.
@@
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . In the modular flavor symmetry framework, the superpotential is tightly constrained by the modular symmetry, with the relevant Yukawa couplings being modular forms of a certain level $N$, which are functions of a modulus field. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been used in various studies to explain the low energy flavor structures of the Standard Model. Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl . Moreover, modular flavor symmetries can be combined with the GUT framework to further enhance the predictive power of the models, such as combining $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
@@
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking. It can also be the only source of CP symmetry breaking Novichkov:2019sqv ; Feruglio:2021dte ; Feruglio:2024ytl . The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms (thus the Yukawa couplings) and the fermion mass matrices are fixed once the modular flavour symmetry is fully or partially broken by the VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with a modular-invariant scalar potential in 4d $N=1$ supergravity. This is also crucial for giving nonzero masses to the modulus field, rendering it phenomenological acceptable.
@@
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that are close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain a dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induce a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory. In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action. Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
@@
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem using modular flavor symmetry. In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield. In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ Sec 6 contains our conclusions.
@@ 2 Modular $\Gamma_{N}$ symmetry
The full modular group is defined as the set of linear fractional transformations acting on the complex modulus $\tau$ in the upper half complex plane
@@
This group can be generated by two generators, $S$ and $T$, which act as follows
@@
These generators satisfy the relations
@@
The $S,T$ generators can be represented by the following $SL(2,Z)$ matrices,
@@
Finite modular groups $\Gamma_{N}$ Feruglio:2017spp are defined as $\Gamma_{N}=\overline{\Gamma}/\overline{\Gamma}(N)$, where $\overline{\Gamma}\equiv SL(2,Z)/\{\pm 1\}$ and $\overline{\Gamma}(N)$ is given by
$\overline{\Gamma}(N)=\left\{\begin{array}[]{c}\Gamma(N),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}N\geq 3\\ \Gamma(N)/\{\pm 1\},~{}~{}~{}N=2\end{array}\right.~{}~{}~{},$ (7)
@@
Modular forms of weight $2k$ and level $N$ span a linear space of finite dimension. They can be arranged into some modular multiplets $Y_{\mathbf{r}}^{(2k)}$ that transform as irreducible representation $\mathbf{r}$ of the finite modular group $\Gamma_{N}$ up to automorphy factor Feruglio:2017spp , i.e.
@@ 3 Natural $\mu$-term in modular flavor model
As noted in Bae:2019dgg , to solve the SUSY $\mu$ problem, the appearance of a $\mu$ term in the superpotential should be prohibited by some symmetry. We propose that the superfields $H_{u}$, $H_{d}$ transform non-trivially under modular transformations, thereby prohibiting the ordinary $\mu H_{u}H_{d}$ term by modular symmetry.
@@
where the coefficient for bilinear $H_{u}H_{d}$ term is given by
$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
for $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ to ensure the modular invariance of
@@
Here, $\alpha\sim{\cal O}(1)$ is a dimensionless parameter and $\Lambda_{SUSY}$ is typically of order the SUSY scale. The Dedekind $\eta$-function is a modular form with modular weight $1/2$ under modular transformation up to a phase. Its transformation under $\gamma$ is given by
$\eta(\gamma\tau)=\epsilon_{\eta}(\gamma)(c\tau+d)^{1/2}\eta(\tau)~{},~{}~{}\forall\gamma\in\overline{\Gamma}~{},$ (15)
where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right],$ (16)
@@
It is evident that the Kahler potential for superfields with non-trivial modular weights, such as $H_{u}$ and $H_{d}$ superfields, is non-minimal. Therefore, after the modulus field $\tau$ acquires a VEV through the modulus stabilization mechanism, the $H_{u},H_{d}$ fields need to be normalized properly to get their canonical kinetic forms. This normalization corresponds to the following replacements
$H_{u}\rightarrow H_{u}(2\Im\tau)^{k_{H_{u}}/2}~{},~{}~{}~{}H_{d}\rightarrow H_{d}(2\Im\tau)^{k_{H_{d}}/2}~{}.$ (18)
Thus, when the modulus field is stabilized, the effective $\mu$-parameter is regenerated as
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$. We will show that, the smallness of $\mu_{eff}$ is the consequence of large suppression factors, either originate from some power of deviation parameter that characterizes the deviation of modulus VEV from typical fixed point, or from non-trivial power (determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of a small factor related to the modulus VEV. All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects.
4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight $-3$ superpotential to determine the minimum. The modulus VEV can always be restricted to the fundamental domain
${D}_{H}\equiv\left\{\tau\in\mathcal{H}:-\frac{1}{2}\leq\Re{\tau}<\frac{1}{2},\,|\tau|>1\right\}\cup\left\{\tau\in\mathcal{H}:-\frac{1}{2}<\Re\tau\leq 0,\,|\tau|=1\right\}\,,$ (20)
of the modular group, as any choice of the modulus VEV in the upper-half plane can be uniquely related to a $\tau\in D_{H}$ in the fundamental domain via a modular transformation, which is physically equivalent to such a choice. Discussions on modulus stabilization with single modulus and multiple moduli can be found in Kobayashi:2019xvz ; Kobayashi:2020hoc ; Kobayashi:2020uaj ; Ishiguro:2020tmo ; NPP:2201.02020 ; PLNSR:2304.14437 ; King:2023snq ; King:2024ssx ; Ding:2024neh ; Higaki:2024pql ; Higaki:2024jdk ; Kobayashi:2023spx ; Funakoshi:2024yxg .
To ensure the modular invariance and Kahler transformation invariance of